Given Bayes formula P(R I8)PE IA) P(A) P(B) Compun tae fellowing P(AIB
Answer #1

P(\overline{A}|\overline{B})=\frac{P(\overline{B}|\overline{A})P(\overline{A})}{P(\overline{B})}

  =\frac{\frac{P(\overline{B}\cap \overline{A}}{P(\overline{A})})P(\overline{A})}{P(\overline{B})}

  =\frac{P(\overline{A}\cap \overline{B})}{P(\overline{B})}

  =\frac{1-P(A\cup B)}{1-P(B)}